OpenStudy (anonymous):

1. Add or subtract terms whenever possible. -3 square root of 162 + 2 square root of 200 - 9 square root of eight 2. Simplify using properties of exponents. 16x^2/3/4x^1/2

OpenStudy (anonymous):

First one looks like this, right?\[-3\sqrt{162}+2\sqrt{200}-9\sqrt{8}\]

OpenStudy (anonymous):

yes. @jabberwock